# Disentanglement in Implicit Causal Models via Switch Variable

### v1

@@ 6 Conclusion
A2.1 $\mathcal{M}\sim_{r}\mathcal{M^{\prime}}\Rightarrow p^{\mathcal{X}}_{\mathcal{M}}(x,\tilde{x})=p^{\mathcal{X}}_{\mathcal{M^{\prime}}}(x,\tilde{x})$
This direction is fairly straightforward. According to Definition 3.6, the fact that $M\sim_{r}M^{\prime}$ implies that $\phi_{E}$ is measure preserving. Therefore, $p^{\mathcal{E}}_{\mathcal{M^{\prime}}}(e^{\prime},\tilde{e^{\prime}})=(\phi_{\mathcal{E}})_{*}p^{\mathcal{E}}_{\mathcal{M}}(e,\tilde{e})$. Furthermore, considering that ancestry is preserved, $\phi_{\mathcal{Z}}$ is measure preserving, and that causal variables are obtained from their ancestral exogenous variables in implicit models, we have $p^{\mathcal{Z}}_{\mathcal{M^{\prime}}}(z^{\prime},\tilde{z^{\prime}})=(\phi_{\mathcal{Z}})_{*}p^{\mathcal{Z}}_{\mathcal{M}}(z,\tilde{z})$. Since models are trained to maximize the log likelihood of $p(x,\tilde{x},\tilde{x}-x)$ and the latent spaces in $M$ and $M^{\prime}$ have the same distribution, the decoders should yield the same observational distributions $p^{\mathcal{X}}_{\mathcal{M}}(x,\tilde{x})=p^{\mathcal{X}}_{\mathcal{M^{\prime}}}(x,\tilde{x})$.
A2.2 $p^{\mathcal{X}}_{\mathcal{M}}(x,\tilde{x})=p^{\mathcal{X}}_{\mathcal{M^{\prime}}}(x,\tilde{x})\Rightarrow\mathcal{M}\sim_{r}\mathcal{M^{\prime}}$
@@
$\forall I\in\mathcal{I}:\>p_{\mathcal{M}}^{\mathcal{Z}}(z,\tilde{z}|I)=p_{\mathcal{M^{\prime}}}^{\mathcal{Z}}(z,\tilde{z}|I)$
@@
Step 2: Component-wise $\phi_{\mathcal{Z}}$ Assuming $V$ is observed or $g$ is linear, we can observe how soft intervention changes $f\rightarrow\tilde{f}$. The causal mechanism switch variable $V$ can be seen as the new and possibly altered set of the parents of causal variables which is why we choose $p(v)$ to be same as $p(e)$ which in our experiments are standard Gaussian. Furthermore, $h$ is used to model the change in effect of the parents due to soft intervention. Each causal variable $Z_{i}$ has its own $h_{i}$ as the soft intervention effect on causal mechanisms may be different for each causal variable. By modeling soft interventions this way, we can disentangle the effect of soft intervention in higher level representations:
  • Observed changes due to intervention effect of causal mechanisms and parents.
  • Observed changes due to intervention effect on causal variable itself which is integrated in $\tilde{E_{i}}$.
By using $V$ as an auxiliary variable we can model the changes in parent set and causal mechanisms. Since we also use single encoder and decoder for $e$ and $\tilde{e}$, the model already sees the effect of $e_{/i}$ on the intervened causal variable and observed variables, hence, the interventional data will be informative about $\tilde{e_{i}}$. Let’s denote all $\mathcal{Z}$ except $\mathcal{Z}_{i}$ as $\mathcal{Z}_{o}$. The connection between the causal variables in $\mathcal{M}$ and $\mathcal{M^{\prime}}$ is shown in Figure 2. Assuming that $\tilde{\mathcal{Z}_{i}}=\tilde{\mathcal{Z}^{\prime}_{i}}=\mathbb{R}$ and Lemma 2 in Brehmer et al. [2022], it can be concluded that the transformation $\tilde{\mathcal{Z}_{i}}\times\mathcal{Z}\rightarrow\tilde{\mathcal{Z}^{\prime}_{i}}$ is constant in $\mathcal{Z}$. Using the diagram in left hand side of the equality in Figure 2(b), it can be further concluded that $\phi_{\mathcal{Z}_{i}}$ is constant in $\tilde{\mathcal{Z}_{o}}$. Therefore, $\phi_{\mathcal{Z}_{i}}$ is a component-wise transformation.
@@
(a) Pre-Epic-Kitchens
(b) Pre-Epic-Kitchens
(c) Pre-Epic-Kitchens
(d) Pre-Epic-Kitchens
(a) Post: Valve-locked
(b) Post: Bread-Inserted
(c) Post: Clothes-Gathered
(d) Post: Juice-Poured
(e) Pre-ProcTHOR
(f) Pre-ProcTHOR
(g) Pre-ProcTHOR
(h) Pre-ProcTHOR
(e) Post: Cabinet-Open
(f) Post: Box-Open
(g) Post: TV-Broken
(h) Post: TV-On
Figure A5: In the Causal-Triplet dataset Liu et al. [2023], visual representations capture both pre and post-intervention scenarios. The first two rows showcase data samples from Epic-Kitchens, while the third and fourth rows feature samples from ProcTHOR. Each image in the post-intervention condition is accompanied by labels specifying the corresponding action and intervened object. In the images in the first two rows, the agent is performing an action on an object but the camera angle has also changed. So we can say that for example the distribution of causal variables conditioned on the camera angle has been changed due to soft intervention.
@@
In a causal model, an intervention refers to a deliberate action taken to manipulate or change one or more variables in order to observe its impact on other variables within the causal model. Interventions help to study how changes in one variable directly cause changes in another, thereby revealing causal relationships. Based on the levels of control and manipulation in a causal intervention, we can have soft vs. hard interventions. A hard intervention involves directly manipulating the variables of interest in a controlled manner such as Randomized Controlled Trials (RCTs). In other words, a hard intervention sets the value of a causal variable $Z$ to a certain value denoted as $do(Z=z)$ Pearl et al. [2016]. On the other hand, soft intervention involves more subtle or less controlled manipulation of variables and changes the conditional distribution of the causal variable $p(Z|Z_{pa})\rightarrow\tilde{p}(Z|Z_{pa})$ which can be modeled as $\tilde{z_{i}}=\tilde{f}_{i}(z_{pa_{i}},\tilde{e}_{i})$ Correa and Bareinboim [2020]. Looking at interventions from a graphical standpoint, a hard intervention entails that the intervened node is solely impacted by the intervention itself, with no influence coming from its ancestral nodes. Conversely, in the context of a soft intervention, the representation of the intervened node can be influenced not only by the intervention but also by its parent nodes. As an example, suppose we are trying to understand the causal relationship between different types of diets and weight loss. The soft intervention in this scenario could be a switch from a regular diet to a low-carb diet. Switching to a low-carb diet is a voluntary choice made by the individual and there are no external forces or regulations compelling them to make this change (non-coercive). The intervention involves a modification of the individual’s diet rather than a complete disruption since they are adjusting the proportion of macronutrients (fats, proteins, and carbs) they consume, which is less disruptive than a radical change in eating habits (gradual modification). The individual has autonomy to choose and tailor their diet according to their preferences and health goals so they are empowered to make informed decisions about their dietary choices (behavioural empowerment). Conversely, if the government or an authority were to intervene and enforce a mandatory low-carb diet through legal means, this would constitute a hard intervention. In this scenario, regulations would be implemented, prohibiting the consumption of specific carbohydrate-containing foods. Regulatory agencies would be established to oversee and ensure adherence to the low-carb diet mandate, taking actions such as removing prohibited foods from the market, restricting their import and production, and so on. Individuals caught consuming banned foods would be subject to fines, legal repercussions, or other penalties.
@@
This section contains additional details about SoftCD design architectures, datasets, and experiments settings.
A4.1 Datasets
The Causal-Triplet datasets are consisted of images containing objects in which an action is manipulating the objects shown in Figure A5. Examples of actions and objects in these datasets are given in Table A2.
@@
Based on the actions and objects, we treat our causal variables as attributes of objects which can be changed by actions. Therefore, actions in these datasets are considered as interventions. Assume that $z_{1}$ corresponds to the attributes of an object, e.g. a door, the target of opening or closing (action’s target) is $z_{1}$. We use actions’ labels in these datasets to detect the targets of interventions to determine which causal variable has been intervened upon. Note that informing the model about the target of intervention is not same as informing about the action itself (See Table 3).
A4.2 Architecture Design
@@
Based on the SoftCD architecture depicted in Figure 1, we devised a solution function comprising three fully connected networks. These networks consist of two layers each, with 64 hidden units per layer and ReLU activation functions. For our synthetic dataset experiments, we employed a fully connected auto-encoder model, while for the realistic dataset, we utilized ResNet-based networks.
Appendix A5 Results
In our synthetic dataset experiments, SoftCD achieved the highest average performance with a 4-variable causal model. We employed various data generation seeds, resulting in diverse adjacency matrices, reflecting the influence of soft interventions on causal mechanisms and parental impact.
A5.1 Scalability
@@
While our primary research objective centered on addressing identifiability challenges in implicit causal models under soft interventions, we also conducted an investigation into the scalability of our proposed model. To comprehensively assess its performance, we designed experiments covering a range of causal graphs, featuring 5 to 10 variables, with 10 different seeds for each variable, following a similar experimental setup as our 4-variable causal graph experiments. The outcomes of these experiments, comparing SoftCD and ILCM, are presented in Figure A6. By increasing the number of variables in the graph, confounding factors and ambiguities of causal relations increase as well. Consequently, more supervision on $\mathcal{V}$ is required to better separate the effect of causal variables themselves on the observed variables.

### v2

@@ 6 Conclusion
A1.1 $\mathcal{M}\sim_{r}\mathcal{M^{\prime}}\Rightarrow p^{\mathcal{X}}_{\mathcal{M}}(x,\tilde{x})=p^{\mathcal{X}}_{\mathcal{M^{\prime}}}(x,\tilde{x})$
This direction is fairly straightforward. According to Definition 3.4, the fact that $M\sim_{r}M^{\prime}$ implies that $\phi_{\mathcal{E}}$ is measure preserving. Therefore, $p^{\mathcal{E}}_{\mathcal{M^{\prime}}}(e^{\prime},\tilde{e^{\prime}})=(\phi_{\mathcal{E}})_{*}p^{\mathcal{E}}_{\mathcal{M}}(e,\tilde{e})$. Furthermore, considering that ancestry is preserved, $\phi_{\mathcal{Z}}$ is measure preserving, and that causal variables are obtained from their ancestral exogenous variables in implicit models, we have $p^{\mathcal{Z}}_{\mathcal{M^{\prime}}}(z^{\prime},\tilde{z^{\prime}})=(\phi_{\mathcal{Z}})_{*}p^{\mathcal{Z}}_{\mathcal{M}}(z,\tilde{z})$. Since models are trained to maximize the log likelihood of $p(x,\tilde{x},\tilde{x}-x)$ and the latent spaces in $M$ and $M^{\prime}$ have the same distribution, the decoders should yield the same observational distributions $p^{\mathcal{X}}_{\mathcal{M}}(x,\tilde{x})=p^{\mathcal{X}}_{\mathcal{M^{\prime}}}(x,\tilde{x})$.
A1.2 $p^{\mathcal{X}}_{\mathcal{M}}(x,\tilde{x})=p^{\mathcal{X}}_{\mathcal{M^{\prime}}}(x,\tilde{x})\Rightarrow\mathcal{M}\sim_{r}\mathcal{M^{\prime}}$
@@
$\forall I\in\mathcal{I}:\>p_{\mathcal{M}}^{\mathcal{Z}}(z,\tilde{z}|I)=p_{\mathcal{M^{\prime}}}^{\mathcal{Z}}(z,\tilde{z}|I)$ (9)
@@
Step 2: Component-wise $\phi_{\mathcal{Z}}$
According to Lemma A1.5, in order to prove that $\phi_{\mathcal{Z}}$ is a component-wise transformation, we need to prove that $\tilde{\mathcal{E}^{\prime}_{i}}$ and $\tilde{\mathcal{E}^{\prime}_{j}}$ are independent $\forall i,j,i\neq j$. In implicit modeling we do not know the parents of each causal variable, hence, we assume the distribution of $\tilde{\mathcal{Z}^{\prime}_{i}}$ to be conditioned only on $\mathcal{E}^{\prime}_{i}$ as in Equation 5 since $\mathcal{E}^{\prime}_{i}$ is a known parent of $\tilde{\mathcal{Z}^{\prime}_{i}}$. The mean of a conditional distribution can be calculated as:
$E[\tilde{z}^{\prime}_{i}|e^{\prime}_{i}]=\mu_{\tilde{z}^{\prime}_{i}}+\rho\frac{\sigma_{\tilde{z}^{\prime}_{i}}}{\sigma_{e^{\prime}_{i}}}(e^{\prime}_{i}-\mu_{e^{\prime}_{i}})$ (10)
where $\rho$ and $\sigma$ are the correlation coefficient and variance of the random variables, respectively. On the other hand, we model $\tilde{\mathcal{Z}^{\prime}_{i}}$ using switch mechanisms as:
$\tilde{z}^{\prime}_{i}=s_{i}(\tilde{e}^{\prime}_{i};e^{\prime}_{/i},h(v^{\prime}))$
By using Taylor’s expansion we can write above equation as:
$s_{i}(\tilde{e}^{\prime}_{i};e^{\prime}_{/i},h_{i}(v^{\prime}))=s_{i}(\tilde{e}^{\prime}_{i};e^{\prime}_{/i},h_{i}(v^{\prime}_{0}))++\sum_{n=1}^{\infty}\frac{1}{n!}\left(\frac{\partial^{n}s_{i}}{\partial h_{i}^{n}}\bigg{|}_{h_{i}=h_{i}(v^{\prime}_{0})}(h_{i}(v^{\prime})-h_{i}(v^{\prime}_{0}))^{n}\right)$
$=s_{i}(\tilde{e}^{\prime}_{i};e^{\prime}_{/i},h_{i}(v^{\prime}_{0}))+R_{i}$
Furthermore, we assume separable dependence such that:
$\exists v^{\prime}_{0}\text{ such that }\forall i\quad s_{i}(\tilde{e}^{\prime}_{i};e^{\prime}_{/i},h_{i}(v^{\prime}_{0}))=s_{i}(\tilde{e}^{\prime}_{i};e^{\prime}_{/i})$
An example of such a scenario could be in location-scale noise models, where a soft intervention changes the location parameter of the model as:
$s_{i}(e^{\prime}_{i};e^{\prime}_{/i})=e^{\prime}_{i}+loc(e^{\prime}_{/i})\rightarrow\tilde{s_{i}}(\tilde{e}^{\prime}_{i};e^{\prime}_{/i})=s_{i}(\tilde{e}^{\prime}_{i};e^{\prime}_{/i},h_{i}(v^{\prime}))$
$=\tilde{e}^{\prime}_{i}+loc(e^{\prime}_{/i})+h_{i}(v^{\prime})=\tilde{e}^{\prime}_{i}+loc(e^{\prime}_{/i})+v^{\prime 2}+v^{\prime}$
In this example, for $v_{0}^{\prime}=0$, $s_{i}(\tilde{e}^{\prime}_{i};e^{\prime}_{/i},h_{i}(v^{\prime}_{0}))=s_{i}(\tilde{e}^{\prime}_{i};e^{\prime}_{/i})$.
Consequently, we can write the following equality from Equation 10:
$E[\tilde{\mathcal{Z}^{\prime}_{i}}|e^{\prime}_{i}]=E[s_{i}(\tilde{\mathcal{E}^{\prime}_{i}};\mathcal{E}^{\prime}_{/i})+R_{i}|e^{\prime}_{i}]=\mu_{\tilde{\mathcal{Z}^{\prime}_{i}}}+\rho\frac{\sigma_{\tilde{\mathcal{Z}^{\prime}_{i}}}}{\sigma_{\mathcal{E}^{\prime}_{i}}}(e^{\prime}_{i}-\mu_{\mathcal{E}^{\prime}_{i}})$
By taking the partial derivative of both side with respect to $\tilde{\mathcal{E}^{\prime}_{j}}$ we have:
$\forall j\neq i\quad E[\frac{\partial s_{i}(\tilde{\mathcal{E}}^{\prime}_{i};\mathcal{E}^{\prime}_{/i})}{\partial\tilde{\mathcal{E}}^{\prime}_{i}}\cdot\frac{\partial\tilde{\mathcal{E}}^{\prime}_{i}}{\partial\tilde{\mathcal{E}^{\prime}_{j}}}+\frac{\partial s_{i}(\tilde{\mathcal{E}}^{\prime}_{i};\mathcal{E}^{\prime}_{/i})}{\partial\mathcal{E}^{\prime}_{/i}}\cdot\frac{\partial\mathcal{E}^{\prime}_{/i}}{\partial\tilde{\mathcal{E}^{\prime}_{j}}}+\frac{\partial R_{i}}{\partial\tilde{\mathcal{E}^{\prime}_{j}}}|e^{\prime}_{i}]=0$
If we did not have the causal mechanism switch variable $(h_{i}(\mathcal{V}^{\prime}))$, the equation above would only hold if $s_{i}$ was constant in parents, which is not the case due to the presence of soft interventions, or if $\frac{\partial s_{i}(\tilde{\mathcal{E}}^{\prime}_{i};\mathcal{E}^{\prime}_{/i})}{\partial\tilde{\mathcal{E}}^{\prime}_{i}}\cdot\frac{\partial\tilde{\mathcal{E}}^{\prime}_{i}}{\partial\tilde{\mathcal{E}^{\prime}_{j}}}=-\frac{\partial s_{i}(\tilde{\mathcal{E}}^{\prime}_{i};\mathcal{E}^{\prime}_{/i})}{\partial\mathcal{E}^{\prime}_{/i}}\cdot\frac{\partial\mathcal{E}^{\prime}_{/i}}{\partial\tilde{\mathcal{E}^{\prime}_{j}}}$. The latter scenario would imply that $\frac{\partial\tilde{\mathcal{E}^{\prime}_{i}}}{\partial\tilde{\mathcal{E}^{\prime}_{j}}}\neq 0$, hence, $\tilde{\mathcal{E}^{\prime}_{i}}\not\perp\!\!\!\perp\tilde{\mathcal{E}^{\prime}_{j}}$. However, by introducing the causal mechanism switch variable $\mathcal{V}$ and assuming it is observed, we can account for the effects of soft interventions through $h_{i}(\mathcal{V}^{\prime})$. In this case, $\frac{\partial\tilde{\mathcal{E}^{\prime}_{i}}}{\partial\tilde{\mathcal{E}^{\prime}_{j}}}=0$ as exogenous variables are commonly assumed to be independent in practice. Consequently:
$\forall i,j\quad\tilde{\mathcal{E}^{\prime}_{i}}\perp\!\!\!\perp\tilde{\mathcal{E}^{\prime}_{j}}$
$\rightarrow\forall i,j\quad p(\tilde{\mathcal{Z}^{\prime}_{i}},\tilde{\mathcal{Z}^{\prime}_{j}}|\tilde{\mathcal{Z}_{i}},\tilde{\mathcal{Z}_{j}})=p(\tilde{\mathcal{Z}^{\prime}_{i}}|\tilde{\mathcal{Z}_{i}})p(\tilde{\mathcal{Z}^{\prime}_{j}}|\tilde{\mathcal{Z}_{j}})$
[…]
@@
(a) Pre-Epic-Kitchens (b) Pre-Epic-Kitchens (c) Pre-Epic-Kitchens (d) Pre-Epic-Kitchens
Refer to caption Refer to caption Refer to caption Refer to caption
(a) Post: Valve-locked (b) Post: Bread-Inserted (c) Post: Clothes-Gathered (d) Post: Juice-Poured
Refer to caption Refer to caption Refer to caption Refer to caption
(e) Pre-ProcTHOR (f) Pre-ProcTHOR (g) Pre-ProcTHOR (h) Pre-ProcTHOR
Refer to caption Refer to caption Refer to caption Refer to caption
(e) Post: Cabinet-Open (f) Post: Box-Open (g) Post: TV-Broken (h) Post: TV-On
Refer to caption Refer to caption Refer to caption Refer to caption
Figure A4: In the Causal-Triplet dataset [19], visual representations capture both pre and post-intervention scenarios. The first two rows showcase data samples from Epic-Kitchens, while the third and fourth rows feature samples from ProcTHOR. Each image in the post-intervention condition is accompanied by labels specifying the corresponding action and intervened object. In the images in the first two rows, the agent is performing an action on an object but the camera angle has also changed. So we can say that for example the distribution of causal variables conditioned on the camera angle has been changed due to soft intervention.
@@
In a causal model, an intervention refers to a deliberate action taken to manipulate or change one or more variables in order to observe its impact on other variables within the causal model. Interventions help to study how changes in one variable directly cause changes in another, thereby revealing causal relationships.
Figure A5: Causal graph models in the presence of Hard (a) and Soft (b) interventions. There are no connections from parents to $\tilde{Z}_{i}$ in hard interventions (a). Whereas, parents are connected to $\tilde{Z}_{i}$ in soft interventions (b).Let’s consider an implicit model and use $/i$ to denote all variables except variable $i$. The major difference of soft intervention (b) with hard intervention (a) is that $\tilde{\mathcal{Z}_{i}}$ is no longer disconnected from its parents and its causal mechanism $\tilde{s}_{i}$ is affected by the intervention. Thus, with a hard intervention, we know the post-intervention parents of a node $\tilde{\mathcal{Z}}_{i}$ (there are none), whereas with soft interventions, the parents themselves may not change.
Based on the levels of control and manipulation in a causal intervention, we can have soft vs. hard interventions. A hard intervention involves directly manipulating the variables of interest in a controlled manner such as Randomized Controlled Trials (RCTs). In other words, a hard intervention sets the value of a causal variable $Z$ to a certain value denoted as $do(Z=z)$ [24].
On the other hand, soft intervention involves more subtle or less controlled manipulation of variables and changes the conditional distribution of the causal variable $p(Z|Z_{pa})\rightarrow\tilde{p}(Z|Z_{pa})$ which can be modeled as $\tilde{z_{i}}=\tilde{f}_{i}(z_{pa_{i}},\tilde{e}_{i})$ [7].
Looking at interventions from a graphical standpoint, a hard intervention entails that the intervened node is solely impacted by the intervention itself, with no influence coming from its ancestral nodes. Conversely, in the context of a soft intervention, the representation of the intervened node can be influenced not only by the intervention but also by its parent nodes.
As an example, suppose we are trying to understand the causal relationship between different types of diets and weight loss. The soft intervention in this scenario could be a switch from a regular diet to a low-carb diet. Switching to a low-carb diet is a voluntary choice made by the individual and there are no external forces or regulations compelling them to make this change (non-coercive).
The intervention involves a modification of the individual’s diet rather than a complete disruption since they are adjusting the proportion of macronutrients (fats, proteins, and carbs) they consume, which is less disruptive than a radical change in eating habits (gradual modification). The individual has autonomy to choose and tailor their diet according to their preferences and health goals so they are empowered to make informed decisions about their dietary choices (behavioural empowerment).
Conversely, if the government or an authority were to intervene and enforce a mandatory low-carb diet through legal means, this would constitute a hard intervention. In this scenario, regulations would be implemented, prohibiting the consumption of specific carbohydrate-containing foods. Regulatory agencies would be established to oversee and ensure adherence to the low-carb diet mandate, taking actions such as removing prohibited foods from the market, restricting their import and production, and so on. Individuals caught consuming banned foods would be subject to fines, legal repercussions, or other penalties.
@@
This section contains additional details about ICRL-SM design architectures, datasets, and experiments settings.
A3.1 Datasets
A3.1.1 Synthetic
[…]
where the $loc$ and $scale$ networks are changed in post intervention. The pre-intervention $loc$ and post-intervention $\widetilde{loc}$ network weights are initialized with samples drawn from $\mathcal{N}(0,1)$ and $\mathcal{N}(3,1)$, respectively. For ablation studies, we change the mean of these Normal distributions. The $scale$ is constant 1 for both pre-intervention and post-intervention samples. Both $e_{i}$ and $\tilde{e_{i}}$ are sampled from a standard Gaussian. The causal variables are mapped to the data space through a randomly sampled $SO(n)$ rotation. For each dataset, we generate 100,000 training samples, 10,000 validation samples, and 10,000 test samples.
A3.1.2 Causal-Triplet
The Causal-Triplet datasets are consisted of images containing objects in which an action is manipulating the objects shown in Figure A4. Examples of actions and objects in these datasets are given in Table A1 and A2.
@@
Based on the actions and objects, we treat our causal variables as attributes of objects which can be changed by actions. Therefore, actions in these datasets are considered as interventions. Assume that $z_{1}$ corresponds to the attributes of an object, e.g. a door, the target of opening or closing (action’s target) is $z_{1}$.
We use actions’ labels in these datasets to detect the targets of interventions to determine which causal variable has been intervened upon. Note that informing the model about the target of intervention is not same as informing about the action itself (See Table 3). We use 5000 images of these datasets to train all models.
A3.2 Architecture Design
@@
Based on the ICRL-SM architecture depicted in Figure 2(a), we devised a location-scale solution function (Equation 6) in which the $loc_{i}$ and $scale_{i}$, and $h_{i}$ networks each comprise of fully connected networks. These networks consist of two layers each, with 64 hidden units per layer and ReLU activation functions. The encoder and decoder parameters for latents $\mathcal{E}$ and $\tilde{\mathcal{E}}$ are shared and we use a separate encoder and decoder with the same architecture for the latent $\mathcal{V}$. For our synthetic dataset experiments, the encoder and decoder are consisted of fully connected networks with 2 hidden layers and 64 units in each hidden layer. For the Causal-Triplet datasets, we utilized ResNet-based networks. The same encoder and decoder architectures are used for all baseline models in the experiments. ResNet50 encoder, ResNet50 decoder, and classifiers with 1 hidden layer and 64 hidden units are used for predicting actions and objects for experiments in Table 4 and Table 3. ResNet18 encoder, ResNet18 decoder, and classifiers with 2 hidden layer and 2 hidden units are used for predicting actions and objects for experiments in Table A4 and Table A3.
A3.3 Training
To enforce the condition described in Equation 5 for $i\notin\mathcal{I}$, we assign the post-intervention exogenous variables the same value as the pre-intervention exogenous variables. In mathematical terms, this translates to $\forall i\notin\mathcal{I}$, we set $\tilde{e}_{i}=e_{i}$.
In our experiments, we do not pretrain the networks, however, for the baseline models we follow the training procedure in [3]. We also use consistency in our experiments to ensure that the encoder and decoder are inverse of each other. Consistency regularizer is used as $\sum_{i}E_{\hat{x}\sim p(\hat{x}|e),x\sim p(x)}[(x-\hat{x})^{2}]$ where $\hat{x}$ are the reconstructed samples.
For optimization, Adam optimizer is used with default hyperparamters. In the synthetic experiments, learning rate changes from $43e-$ 4 to $81e-$ 8 with a cosine scheduler. In the Causal-Triplet experiments in Table 4 and Table 3 learning rate changes from $0.002$ to $81e-$ 8 with a cosine scheduler. For Table A4 and Table A3 experiments earning rate changes from $0.0001$ to $81e-$ 8 with a cosine scheduler. In all experiments the batch size is set to 64. In the main Causal-Triplet experiments we train the models for 400 epochs, in the appendix Causal-Triplet experiments we train the models for 2000 epochs, and in the synthetic experiments we train the models for 100 epochs. In the appendix experiments, the graph parameters for explicit models are frozen after 1000 epochs.
All models are trained using Nvidia GeForce RTX4090 GPUs. Each of the Causal-Triplet experiments takes 3-8 hours to train the models and each of the synthetic experiments takes 2-3 hours to train the models.
We save the models’ weights with best validation loss and evaluate them using those weights with test data.
Appendix A4 Ablation study
A4.1 Scalability
@@
While our primary research objective centered on addressing identifiability challenges in implicit causal models under soft interventions, we also conducted an investigation into the scalability of our proposed model. To comprehensively assess its performance, we designed experiments covering a range of causal graphs, featuring 5 to 10 variables, with 10 different seeds for each variable, following a similar experimental setup as our 4-variable causal graph experiments. The outcomes of these experiments, comparing ICRL-SM and ILCM, are presented in Figure A6. By increasing the number of variables in the graph, confounding factors and ambiguities of causal relations increase as well. Consequently, more supervision on $\mathcal{V}$ is required to better separate the effect of causal variables themselves on the observed variables.
@@
A4.2 Backbone model
We trained the models using a simpler backbone model, ResNet18, to see how it affects performance. The input image resolution is $64\times 64$ and we use the intervened causal variables to predict action and object classes. The results are shown in Table A4 and A3. It can be seen from the results that the proposed method outperforms other explicit and implicit models even with a simpler model.
Table A3: Table comparing action and object accuracy across various methods on Causal-Triplet datasets using ResNet18 model.
Epic-Kitchens ProcTHOR
Method Action Accuracy Object Accuracy Action Accuracy Object Accuracy
$\beta-VAE$ [11] 0.15 0.04 0.20 0.36
$d-VAE$ [21] 0.16 0.02 0.15 0.38
ILCM [3] 0.19 0.04 0.15 0.42
ICRL-SM (ours) 0.35 0.04 0.40 0.69
Table A4: Action and object accuracy of three explicit models are compared with ICRL-SM. Experiments are conducted applying image with resolution of $R_{64}$ as the input to the Resnet18 encoder with the intervened casual variable ($z_{i}$).
Datasets Methods Action Accuracy Object Accuracy
Epic-Kitchens ENCO [16] 0.14 0.03
DDS [5] 0.16 0.05
Fixed-order 0.14 0.05
ICRL-SM (ours) 0.35 0.04
ProcTHOR ENCO [16] 0.16 0.28
DDS [5] 0.34 0.35
Fixed-order 0.34 0.38
ICRL-SM (ours) 0.40 0.69